OpenStudy (anonymous):

Use natural logarithms to solve the equation. Round to the nearest thousandth. 6e4x - 2 = 3 A. –0.448 B. 0.327 C. 0.067 D. –0.046

OpenStudy (anonymous):

\[6e ^{4x-2}=3\] \[e ^{4x-2}=\frac{ 3 }{6 }=\frac{ 1 }{ 2 }\] \[\ln e ^{4x-2}=\ln \frac{ 1 }{ 2 }=\ln 1-\ln 2 =0-\ln 2\] \[\left( 4x-2 \right)\ln e=-\ln 2\] \[4x-2=-\ln 2 \left(\because \ln e=1 \right)\] \[4x=2-\ln 2\] \[x=\frac{ 1 }{2 }-\frac{ 1 }{ 4 }\ln 2\] after substituting the value of log2 from the calculator or log tables you can find the value of x

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

the answer is A